he linear momentum p of a particle is given as a function of time t as p = at ^2 +Bt + c the dimensioin of constant B are

Grade:11

2 Answers

Shailendra Kumar Sharma
6 years ago
P=At2 + Bt +CDimension of P = mvi.e. MLT^(-1) And now dimension of Bt and C must be same as PSo dimension of C is MLT^(-1) Dimension of B will be MLT^(-2)
jahska
4 years ago
Given,
P=At^2+Bt+C, where A,B and C are constants.
Also,linear momentum(P)=MLT^-1
So, MLT^-1=At^2+Bt+C, C=MLT^-2 [since it is a constant]
therefore,MLT^-1=Bt
                           = MLT^-1/T^1=MLT^-2, Hope it was helpful  :)